i need some help find the slope and reduce p= (3/4, 1 1/4) q=(-1/2,-1) find the slope of pq
Found 2 solutions by Theo, MathTherapy:
Answer by Theo(13342)

let p = (x1, y1)
let q = (x2, y2)

you get:

(x1,y1) = (3/4, 1+1/4)
(x2,y2) = (-1/2, -1)

since 1 + 1/4 is equal to 5/4, this becomes:

(x1,y1) = (3/4, 5/4)
(x2,y2) = (-1/2, -1)

slope is equal to (y2 - y1) / (x2 - x1)
that becomes (-1 - 5/4) / (-1/2 - 3/4)
-1 is equal to -4/4
-1/2 is equal to -2/4
that makes slope equal to (-4/4 - 5/4) / (-2/4 - 3/4)
combine like terms to get slope equal to (-9/4) / (-5/4)
that makes slope equal to -9/4 * -4/5.
that makes slope equal to 9/5.

your slope is equal to 9/5.



Answer by MathTherapy(10552)

i need some help find the slope and reduce p= (3/4, 1 1/4) q=(-1/2,-1) find the slope of pq